

********************************************************************
UARS WEEKLY STATUS REPORT - 29 MAY 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, and
occultation scan data without apparent problems.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on May 20, in the reduced 85 watt power mode.
MLS instrument operations have continued flawlessly. MLS is scheduled to
operate through June 1.
***********
HRDI
***********
HRDI is currently off. Next scheduled turn-on is May 31st.
***********
WINDII
***********
WINDII WEEKLY REPORT FOR THE WEEK ENDING MAY 29, 1998
WINDII was turned on briefly on May 24, but was then turned off for
lack of power.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF and UCSS developer successfully completed the UCSS 10.0 and MSS 4.0
software installation.
The CDHF has received a total of 388 CD-ROM requests as of 29-May-1998.
Two orders were received and shipped.
The CDHF is successfully utilizing the new version of the CSC scheduler
Enhancement (New Expansion Tool) in the Solstice v4.61 processing streams.
The CDHF conducted a survey to determine the usage of the UOAS v4.71
UCSS/UOAS Interactive Services Input/Output screen. The survey was
conducted in regards to PR 2590. Thanks to all for the quick response.
UKMO is experiencing difficulty in transmitting data. They are currently
working on the problem.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-1288C=91L)
60 days HAL (4.80) PROD (L1) (1868R-250C-2F=1616L)
43 days HAL (4.80) PROD (L3) (1868R-44C=1824L)
543 days SOL (4.61) PROD (L1) (2377R-543C=1834L)
76 days SOL (4.61) PROD (L3) (2304R-76C=2228L)
1 day SOL (4.61) PROD (TEST) (1R-1C=0L)
1 day SUS (6.3) PROD (1R-1C=0L)
The CDHF received a request from FDF to begin receiving orbit data on a daily
schedule with a two day time lag. Monday Orbit would be received on Wednesday..
Tuesday Orbit on Thursday etc. We asked that this schedule be put on hold
until all parties involved are notified and in agreement.
CDHF pending software upgrades: IDL 5.1, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 16 May 1998 (Orbit 36484, GMT Day 136) through 22 May
1998 (Orbit 36587, GMT Day 142).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 139/1719 36539 Off
MLS 140/0338 36545 On
PEM 140/2224 36557 ZEPS Off
PEM 141/0204 36559 ZEPS On
The beta angle reached a minimum angle of 0.0 degrees on 05 May 1998
(DOY 125). The beta angle is now increasing toward a maximum angle of 78.0
degrees on 26 May 1998 (DOY 146). Spacecraft full-sun (i.e., no Earth
eclipse) entry began on 21 May 1998 (DOY 141) and will continue until 31
May 1998 (DOY 151).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery
1 remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.8 V end-of-night (EON) load bus voltage.
Battery temperatures are falling (due to high beta angle and low depth-
of-discharge) with a temperature delta between Battery 2 and Battery 3 of
0.8 degrees C. Additional battery performance data for each battery on 16
May 1998 and 22 May 1998 are provided in Appendix D.
The clock error ranged from -9.5 to +12.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = none
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% HALOE 01
% MLS 01
% PEM 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/118 - 2287.497007
Transponder B: 98/120 - 2287.498907
GROUND SYSTEM OPERATION
Two (2) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/134 show the observatory data loss to
be 20 hours, 43 minutes, 22.452 seconds (an increase of 11.296 seconds
since last report period). This is a 0.0356 percent data loss which equals a
99.9644 percent data capture for the mission.
The increase of 11.296 seconds occurred over a two day period:
7 1.664 seconds on Day 131 due to hits in the data, and
7 9.632 seconds on Day 132 due to hits in the data and insufficient
overlap.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/141 show the observatory data loss to
be 20 hours, 43 minutes, 29.652 seconds (an increase of 07.200 seconds
since last report period). This is a 0.0355 percent data loss which equals a
99.9645 percent data capture for the mission.
The increase of 07.200 seconds occurred over a four day period:
7 1.056 seconds on Day 137 due to hits in the data,
7 0.448 seconds on Day 138 due to hits on line,
7 0.512 seconds on Day 140 due to hits in the data, and
7 5.184 seconds on Day 141 due to hits in the data and on line.
FDF reverted back to a daily ephemeris generation and delivery schedule
on 21 May 1998 due to the increasing solar flux. FDF was no longer able to
meet the allowable ephemeris transition tolerance requirement by
generating
ephemeris on a Monday/Wednesday/Friday schedule.
The FOT made a few more modifications to the OBCSTF resident POCC
database this week. These changes were supposed to be the last necessary to
iron out all of the "kinks" in the Database 67 release. However, late this
week, HRDI operations (Dave Gell) expressed interest in operating the
HRDI instrument during spacecraft daylight period only. This new scenario
will require a few additional database modifications. These changes most
likely will be finalized early next week.
UARS completed testing of the 32 kbps Q-channel dumps using the IP net
on 14 May 1998. These tests were performed in parallel with the real-time
operational (NASCOM) system using operational spacecraft data and the
results were mirrored on both systems. However, commanding was
performed only with the real-time system. All 32 kbps dump types were tested -
Ephemeris (system table 33 & 34), Absolute Time Commands (ATC), and OBC
bank dumps. Dump compares were executed automatically by the
Applications Processor (AP) with no errors indicated.
Testing the command side of the I/P Net is on hold. Testing will begin
when MSOCC provides a command loop interface from the POCC through the
I/P Net to the UARS Simulator.
OTHER SIGNIFICANT EVENTS
Continued testing of the upgrade to the OBC Power Monitor (PMON) flight
software. Minor changes to both the flight software and supporting
ground database were made during the week. PMON software modifications and
testing continued. Multiple simulation runs to verify integrity of the
latest modifications, database consistency with the new software, and
testing of new features were executed.
Work continues on the UARS orbit adjust planning scheduled for 05 August
1998. Additional options for an Orbit Adjust are being planned in the
event that the B/D thrusters can not be safely used. Data from the
original anomaly during which the hydrazine lines for the B and D
attitude thrusters were depleted is being prepared for delivery to the
Propulsion Branch (Dennis Asata). The Propulsion Branch stated that they
cannot supply the latch valve driving circuits and that they must be supplied
by some other source to complete the latch valve tests.
UARS continues to assisted TOMS in reviewing the TOMS/EP Transmitter A
failure data.
FOT continues to support CNMOS in their efforts to make necessary
software changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering
effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not
been scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 16 June
1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit
adjust is targeted for 05 August 1998.
PERSONNEL CHANGES
Dimitrios Mantziaras, On-Line Evaluation Engineer transferred to an Off-
Line Engineering position effective 13 May 1998.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 22 May 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
36503 137/0922 98-168 OBC Clk Adj fr 20.6 to 0.00 msec/day
36526 138/2040 98-169 OBC Clk Adj fr 0.00 to 20.6 msec/day
36539 139/1719 98-170 HALOE Power OFF
36545 140/0338 98-171 MLS Power ON
36557 140/2224 98-172 PEM Power OFF ZEPS
36559 141/0204 98-173 PEM Power ON AXIS
36569 141/1735 98-174 OBC Clk Adj fr 20.6 to 41.2 msec/day
36585 142/1920 98-175 OBC Reload RTs 12
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 22 May 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 22 May 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
36503 137/09:29 (17 May 98) from 20.6 to 00.0 msec/day
36526 138/20:44 (18 May 98) from 00.0 to 20.6 msec/day
36569 141/17:35 (21 May 98) from 20.6 to 41.2 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 22 May 1998
16 May 1998 - GMT Day 136 Beta = 45.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.2 0.000 0.0
2 +33.6/ -22.4 +6.11/+4.51 +28.4/-16.0 24.8 1.020 22.3
3 +67.2/ -5.6 +5.15/+3.23 +27.2/-16.8 24.8 1.022 21.8
22 May 1998 - GMT Day 142 Beta = 69.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/ +0.0 18.4 0.000 0.0
2 +44.8/ -11.2 +3.87/+2.29 +20.0/-16.4 24.8 1.004 18.6
3 +67.2/ +0.0 +2.92/+1.34 +19.2/-16.4 24.8 1.012 18.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 22 May 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-021 FOT 36509 137/1724 14170 Generic Late Acquisition
98-021 FOT 36514 138/0141 14170 Generic Late Acquisition
98-024 FOT 36572 141/2323 N/A TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-024 FOT 36572 141/23:27 TAC 4 CRASH
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-021 Generic Late Acq for May 1998 (Open until 31 May 1998).
SEVNTFW21